n a mating between individuals with the genotypes IAi Hh× iiHh, what percentage of the offspring would be expected to have the O blood type?
Answer #1

Answer:

Allele H is dominant over allele h. At least one H allele must be present to express the respective phenotype. In recessive (hh) conditions, all genotype would become O type blood phenotype.

IAi Hh (A type) x (O type) ii Hh ---Parents

iH

ih

IAH

IAi HH (A type)

IAi Hh (A type)

IAh

IAi Hh (A type)

IAi hh (O type)

iH

ii HH (O type)

ii Hh (O type)

ih

ii Hh (O type)

ii hh (O type)

A type = 3/8

O type = 5/8

The percentage of the offspring would be expected to have the O blood type = 5/8 = 62.5%